OpenStudy (anonymous):

@nikato Simplify. 4^2 + 3^3 (2 and 3 are suppose to be above 4 and 3 A. 35 B. 43 C. 97 D. 17

OpenStudy (nikato):

Ok. Lets start with 4^2. Do you know what that is? Remember 4^2 is the same as 4x4

OpenStudy (anonymous):

4x4 = 16 :]

OpenStudy (anonymous):

16 + 16 = 32

OpenStudy (nikato):

Ok. Ur first post is correct. But not ur second, I don't know why you would do 16+16=32. Ok. So u know 4^2=16 Can you now find 3^3?

OpenStudy (anonymous):

Oops I thought like oh I really dont know what I did there lol And 3x3 = 9

OpenStudy (nikato):

No. Remember it's 3 and the 3 is above, which is actually 3x3x3

OpenStudy (anonymous):

o! Thats 27 then.

OpenStudy (nikato):

Correct. In case your still not familiar with exponents which is what we call the little number on the top http://www.mathsisfun.com/exponent.html So now we know 4^2=16 and 3^3=27 So 16+27=?

OpenStudy (anonymous):

= 43. And thank you!

OpenStudy (nikato):

Correct! And your welcome.